PLEASE HELP ASAP THIS IS DUE TOMORROW!
Find the percentage in the shortest way:
a) 25% of 44 books
b) 6 1/4 % of $3,875
c) 16 2/3 % of 96 lamps
d) 87 1/2 % of 540.72
e) 33 1/3 % of 750 boys
f) 66 1/2 % of 240 kg

Answer by Alan3354(69443)

Find the percentage in the shortest way:
a) 25% of 44 books
---
Per cent is per 100.
25% = 25/100 = 1/4
(1/4)*44 = 11 books.
---------
Do the others the same way.
========================================
b) 6 1/4 % of $3,875
c) 16 2/3 % of 96 lamps
d) 87 1/2 % of 540.72
e) 33 1/3 % of 750 boys
f) 66 1/2 % of 240 kg